Circular No. 8157

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gf IN MCG –04-52-26 M. Hamuy, Carnegie Observatories; and J. Maza, University of Chile,

report that a spectrum (range 380–930 nm) of SN 2003gf (cf. IAUC 8156), obtained on June 25.31 UT with the Dupont 2.5-m telescope (+ WFCCD) at Las Campanas, bears resemblance to those of the type-Ic SN 1987M obtained 11–29 days after maximum light (Filippenko 1997, ARAA 35, 309). Superimposed narrow emission Hα emission from the host galaxy yields a recession velocity of 2520 km/s.

SUPERNOVA 2002ic E. Berger and A. M. Soderberg, California Institute of Technology; and

D. A. Frail, National Radio Astronomy Observatory, report an upper limit of 0.1 mJy from an observation of SN 2002ic (IAUC 8019) with the Very Large Array (VLA) at 4.86 GHz on June 16.60 UT; this result is consistent with the dense circumstellar material inferred from Hα (IAUC 8151).

Further to IAUC 8153, C. J. Stockdale, R. A. Sramek, K. W. Weiler, S. D. Van Dyk, and N. Panagia, report a search with the VLA for radio emission from the type-Ia supernova 2002ic (IAUC 8019, 8028, 8151), possibly the first type-Ia event found showing significant interaction between the ejecta and a dense circumstellar medium: “Hamuy et al. (2003,

http://xxx.lanl.gov/abs/astro-ph/?0306270), suggest a similarity to

type-IIn supernovae. Our 3σ upper limit of 1.28 mJy was obtained at 22.485 GHz (wavelength 1.3 cm) on June 17.79 UT (supernova age ≈ 216 days) in ‘A’-configuration at the optical position (IAUC 8019). Comparison with the radio emission from the type-IIn supernova 1988Z (and the very similar SN 1986J) indicates that SN 1988Z, at the distance of SN 2002ic, would have yielded a 5σ detection at this frequency and epoch (cf. Williams et al. 2002, Ap.J. 581, 396). If SN 2002ic were indeed similar to a type-IIn supernova, lower frequencies would be optically thick at this epoch of observation, as is implied by the upper limit obtained by Berger et al. (see above) at 4.86 GHz on the same date. Additional observations are planned.”

NOVA IN M31 K. Hornoch, Lelekovice, Czech Republic, reports the discovery of an ap-

parent nova located at α = 0^{h}43^{m}36^{s}.17, δ = +41^{o}16′39′′.4 (equinox 2000.0), which is 583′′ east and 30′′ north of the center of M31. Approximate R-band magnitudes: June 21.021 UT, [18.3; 25.01, 17.3; 26.02, 16.4.

2003 June 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green